The following transactions and adjusting entries were completed by a paper-packaging company called Gravure Graphics International during 2018 and 2019. The company uses straight-line depreciation for trucks and other vehicles, double-declining-balance depreciation for buildings, and straight-line amortization for patents.

2018
January 2 Paid $101,000 cash to purchase storage shed components.
January 3 Paid $6,000 cash to have the storage shed erected. The storage shed has an estimated life of 10 years and a residual value of $9,000.
April 1 Paid $47,000 cash to purchase a pickup truck for use in the business. The truck has an estimated useful life of five years and a residual value of $7,000.
May 13 Paid $500 cash for minor repairs to the pickup truck's upholstery.
July 1 Paid $14,000 cash to purchase patent rights on a new paper bag manufacturing process. The patent is estimated to have a remaining useful life of five years.
December 31 Recorded depreciation and amortization on the pickup truck, storage shed, and patent.
2019
June 30 Sold the pickup truck for $40,000 cash. (Record the depreciation on the truck prior to recording its disposal.)
December 31 Recorded depreciation on the storage shed. Also determined that the patent was impaired and wrote off its remaining book value (i.e., wrote down the book value to zero).

Required:

Prepare the journal entries required on each of the above dates.

  • 1 Record the purchase of storage shed components for $101,000.

  • 2 Record the payment of $6,000 to have the storage shed erected.

  • 3 Record the purchase of a pickup truck for $47,000.

  • 4 Record the repairs of $500 to the pickup truck.

  • 5 Record the purchase of a patent for $14,000 on a new paper bag manufacturing process.

  • 6 Record the depreciation and amortization expense on the pickup truck, storage shed and patent for the year.

  • 7 Record the depreciation for the truck up to the date of sale.

  • 8 Record the sale of the truck for $40,000 cash.

  • 9 Record the depreciation on the storage shed for the year.

  • 10 Record the Patent Amortization expense for the full year.

  • 11 Record the reversal of the accumulated amortization of patents.

  • 12 Record any impairment loss incurred on the patent.
Answer #1
Journal entries
Event Date General journal Debit Credit
1 Jan 2, 2018 Storage Shed (Building)       101,000
Cash       101,000
(To Record the purchase of storage shed components for $101,000.)
2 Jan 3, 2018 Storage Shed (Building)            6,000
Cash            6,000
(To Record the payment of $6,000 to have the storage shed erected.)
3 Apr 1, 2018 Truck (Equipment)          47,000
Cash          47,000
(To Record the purchase of a pickup truck for $47,000.)
4 May 13, 2018 Repair expense                500
Cash                500
(To Record the repairs of $500 to the pickup truck.)
5 Jul 1, 2018 Patent          14,000
Cash          14,000
(To Record the purchase of a patent for $14,000 on a new paper bag manufacturing process.)
6 Dec 31, 2018 Depreciation expense (9800+6000)          15,800
Amortization expense            1,400
Accumulated depreciation - Storage Shed (Building) (101000+6000=107000) ((107000-9000)/10)            9,800
Accumulated depreciation - Truck (Equipment) ((47000-7000)/5) = 8000 (Apr to Dec = 9 months) (8000*9/12)            6,000
Accumulated Amortization - Patent (14000/5=2800) (Jul to Dec = 6 months) (2800*6/12)            1,400
(To record adjustment entry for Depreciation and amortization expense)
7 Jun 30, 2019 Depreciation expense - Truck (Equipment)            4,000
Accumulated depreciation - Truck (Equipment)            4,000
(To record depreciation for pickup truck.) (Jan to Jun = 6 months) (8000*6/12)
8 Jun 30, 2019 Cash          40,000
Accumulated depreciation - Truck (Equipment) (6000+4000)          10,000
Gain on sale of Truck            3,000
Truck (Equipment)          47,000
(To record sale of the truck .) (Book value = 47000-10000 = 37000) (Gain = 40000-37000)
9 Dec 31, 2019 Depreciation expense            9,800
Accumulated depreciation - Storage Shed (Building)            9,800
(To record depreciation for Storage Shed.)
10 Dec 31, 2019 Amortization expense            2,800
Accumulated Amortization - Patent            2,800
(To record Amortization for Patent.)
11 Dec 31, 2019 Accumulated Amortization - Patent            4,200
Patent            4,200
(To Record the reversal of the accumulated amortization of patents.) (1400+2800)
12 Dec 31, 2019 Impairment expense - patent            9,800
Patent            9,800
(To Record any impairment loss incurred on the patent.) (Book value =14000-4200 )